### v1

# All-DNA associative polymers

Abstract
Associative polymers (APs) with reversible, specific interactions between “sticker” sites exhibit a phase behavior that depends on a delicate balance between distinct contributions controlling the binding. For highly-bonded systems, it is entropy that mostly determines if, on increasing concentration, the network forms progressively or via a first-order transition. With the aim of introducing an experimentally-viable system tailored to test the subtle dependence of the phase behavior on the binding site topology, here we numerically investigate AP polymers made of DNA, where “sticker” sites made by short DNA sequences are interspersed in a flexible backbone of poly-T spacers. Due to their self-complementarity, each binding sequence can associate with another identical sticky sequence. We compare two architectures: one with a single sticker type, $(AA)_{6}$, and one with two distinct alternating types, $(AB)_{6}$. At low temperature, when most of the stickers are involved in a bond, the $(AA)_{6}$ system remains homogeneous, while the $(AB)_{6}$ system exhibits phase separation, driven primarily by entropic factors, mirroring predictions from simpler bead-spring models. Analysis of bond distributions and polymer conformations confirms that the predominantly entropic driving force behind this separation arises from the different topological constraints associated with intra- versus inter-molecular bonding. Our results establish DNA APs as a controllable, realistic platform for studying in the laboratory how the thermodynamics of associative polymer networks depends on the bonding site architecture in a clean and controlled way.
I Introduction
Associative polymers (APs) are a class of macromolecules that contain specific interaction sites, often referred to as “stickers”, distributed along an otherwise inert polymeric backbone [1, 2]. The stickers can reversibly bind to each other, giving rise to a wide range of physical behaviors, including self-assembly, phase separation, and the formation of transient networks [3, 4, 5]. Such dynamic and tunable properties make associative polymers highly relevant both in materials science and in the study of biological systems [6, 7, 8, 9, 10].
In recent years, associative polymers have been extensively employed as model systems to understand the behavior of intrinsically disordered proteins (IDPs) and other biopolymers, which often rely on multivalent, weak interactions to drive phase separation and compartmentalization without membranes [11, 12]. To capture these complex behaviors, associative polymers are typically modeled using coarse-grained representations, where the backbone is treated as a polymer chain and the stickers as discrete sites capable of specific pairwise interactions [1, 2]. In the case of simple models, where chains are flexible, spacers interact through excluded volume only, and each sticker can be involved in no more than one bond, there is theoretical [3], numerical [13, 14, 15], and experimental [7, 16] evidence that no thermodynamic instability is encountered upon increasing concentration. Indeed, in the dilute regime most of the bonds involve stickers belonging to the same chain (intra-molecular bonds), which therefore resembles a single-chain nanoparticle [17]. However, as concentration increases, chains start to interact more strongly with each other, forming inter-molecular connections that are entropically favoured [18, 19], and therefore progressively replace the intra-molecular bonds, thereby giving raise to a network-like structure without any phase separation occurring. However, it has been recently shown that in simple bead-spring models the different entropic contributions that control the overall bonding process can be manipulated by introducing multiple sticker types, with the constraint that only like stickers can bind to each other. Interestingly, as soon as there are at least two sticker types alternating along the backbone, an entropy-driven phase separation appears for a wide range of system parameters [20, 21]. The main contribution dictating the phase behaviour is due to the much larger entropic cost of closing an intra-molecular bond compared to an inter-molecular bond, effectively enhancing the number of inter-molecular contacts, and therefore the overall tendency to phase separate.
To verify that the unexpected results reported in recent numerical studies [20, 21] — based on coarse-grained polymer models — are indeed accurate, it is essential to reproduce them using a realistic polymer system that can be tested experimentally. To this end, we propose a model that, although still investigated numerically, closely mirrors a system that can be realized in the laboratory. Specifically, we represent associative polymers using a carefully designed single-stranded DNA. This is achieved by leveraging the selectivity of short base-pair binding sequences embedded within an otherwise inert poly-thymine (poly-T) strand to encode the desired binding architecture. As a result of this design, each DNA polymer (single strand) consists of several short self-complementary sequences acting as stickers, separated by poly-thymine spacers that provide flexibility and steric separation. This architecture allows precise control over valency and interaction specificity, making it an ideal platform for probing the phase behavior and network formation of associative polymers at the molecular level. In particular, here we design DNA strands that mimic APs decorated with one or two types of stickers in order to provide an experimentally-amenable realisation of the entropic phase separation reported in Refs. [20, 21]. We find that, as for the bead-spring model, DNA APs with a single sticker type does not phase separate, whereas phase separation is observed for two-sticker-type DNA strands. We find that also in the DNA model, despite the presence of electrostatic, excluded volume and bending energy contributions, the separation phenomenon still retains a large favourable entropic term arising from the combinatorial competition between intra- and inter-molecular bonds, as seen in the simple bead-spring model.
II Methods
We simulate two one-component model systems made of single strand DNA in solution that we will call $(AA)_{6}$ and $(AB)_{6}$. Each DNA strand is composed by $N_{\rm nucl}=158$ nucleotides, with a backbone of thymine bases (the spacers) in which are inserted $M=12$ self-complementary regions (the stickers) as shown in Fig. 1, which also reports the full sequences of the strands used throughout this work. In the $(AA)_{6}$ systems the stickers have all the same sequence (TCGA), while in the $(AB)_{6}$ systems the stickers have alternating TGCA and GTAC sequences. The two sequences are self-complementary, so that only stickers of the same type can bind to each other.
Figure 1: (a) The sequences of the (top) $(AA)_{6}$ and (bottom) $(AB)_{6}$ DNA strands. (b)-(c) The sequences of the $A$ and $B$ stickers. (d) The conformation of an $(AA)_{6}$ strand extracted from a simulation, together with two other strands that are bonded to it, coloured in semi-transparent grey. The spacers and stickers of the main strand are depicted in dark grey and green, respectively, while the orange and blue circles highlight an intra-molecular and an inter-molecular bond, respectively. (e) Same as (d) but for an $(AB)_{6}$ strand; in this case there are two types of stickers, which are coloured in green and violet as in (b) and (c).
[…]
We run NVT simulations at several temperatures and densities starting from an initial configuration where $N_{s}=128$ strands are placed randomly in the box (homogeneous simulations). To test for the presence of a phase separation, we also perform direct-coexistence simulations by taking an equilibrium configuration of a homogeneous simulation and elongating the box size along the $x$ direction by a factor of 4, effectively putting it in coexistence with vacuum. We then let the system evolve, monitoring the behaviour of the two initial interfaces.
To identify the gas and liquid regions in the density profile, we fit 10 equilibrium profiles extracted at regular intervals along the trajectory with a smooth analytical function based on hyperbolic tangent transitions, which are well-suited for describing diffuse interfaces in finite systems. Specifically, we model the density profile as:
$\rho(x)=\frac{\rho_{\mathrm{liq}}+\rho_{\mathrm{gas}}}{2}-\frac{\rho_{\mathrm{liq}}-\rho_{\mathrm{gas}}}{2}\left[\tanh\left(\frac{x-x_{0}}{\xi}\right)-\tanh\left(\frac{x-x_{1}}{\xi}\right)\right]\left[\tanh\left(\frac{x-x_{2}}{\xi}\right)-\tanh\left(\frac{x-x_{3}}{\xi}\right)\right]$ (1)
where $\rho_{\text{liq}}$ and $\rho_{\text{gas}}$ are the liquid and gas densities, $\xi$ is the interfacial width, and the positions $x_{0}<x_{1}<x_{2}<x_{3}$ define the boundaries of the interfaces. Specifically, $x<x_{0}$ and $x>x_{3}$ are classified as gas regions, $x_{1}<x<x_{2}$ as liquid, while the segments between $x_{0}$ and $x_{1}$, and between $x_{2}$ and $x_{3}$ correspond to the two interfaces.
During all simulations we keep track of pairs of nucleotides sharing hydrogen bonds (HBs), where two nucleotides are considered bonded if their hydrogen bond interaction energy is lower than $-0.1$ in oxDNA internal units, classifying the bond as intra- or inter-molecular if the two nucleotides belong to the same strand or to different strands, respectively. We then consider two stickers as bonded if their nucleotides are involved in at least two bonds. We thus obtain the total number of stickers involved in intra- and inter-molecular bonds, $N_{\rm intra}$ and $N_{\rm inter}$, and define the fraction of stickers involved in intra- and inter-molecular as $f_{\rm intra}=N_{\rm intra}/N_{\rm max}$ and $f_{\rm inter}=N_{\rm inter}/N_{\rm max}$, where $N_{\rm max}=MN_{s}$ is the total number of stickers. We also define the fraction of bonded stickers as $f_{\rm tot}=(N_{\rm intra}+N_{\rm inter})/N_{\rm max}$.
III Results
Figure 2: The fraction of hydrogen bonds (HBs) between nucleotides belonging to stickers involved in total, intra- or inter-molecular bonds for the (left) $(AA)_{6}$ and (right) $(AB)_{6}$ systems. Each row represents a different density which takes the values, from top to bottom, of $\rho\sigma^{3}=0.57$, $\rho\sigma^{3}=0.31$, and $\rho\sigma^{3}=0.16$, respectively.
We start by characterizing the bonding process in the two systems as a function of temperature and density. Figure 2 shows the temperature-dependence of the fractions of hydrogen-bonded nucleotides belonging to stickers of the $(AA)_{6}$ and $(AB)_{6}$ homogeneous systems, classified according to the type of bond (inter- or intra-molecular), as well as the sum of the two, for three different densities.
These melting curves exhibit the well-known sigmoidal shape typical of DNA pairing, characterized by a density dependent melting temperature which decreases as the density decreases. In almost all simulations, the fraction of bonded nucleotides reaches 90%, signaling the extensive bonding process that takes place in both systems. Consistent with previous results derived for a bead-spring model [20, 21], we observe a striking difference in the fraction of intra-strand bonds between the $(AA)_{6}$ and $(AB)_{6}$ systems. In the $(AB)_{6}$ case the majority of the bonds are formed between sticky sequences of different strands, and the fraction of intra-strand bonds never exceeds 15%. Differently, in the case of the $(AA)_{6}$ system, the fraction of intra bonds is enhanced, becoming dominating at the lowest densities.
We also note on passing that the fraction of intra-molecular bonds nearly always displays a maximum at some intermediate temperature. We posit that this non-monotonic behaviour is due to the finite flexibility of the strands: indeed, as temperature decreases, nucleotides in the spacer regions tend to stack more, making the strands less flexible, thereby effectively hindering the formation of intra-molecular bonds.
To answer the question whether the difference in sequences has indeed a relevant role in controlling the phase behavior of associative polymers in the limit of extensive bonding, we perform direct-coexistence simulations for temperatures lower than $\approx 20^{\circ}$ C, starting from initial configurations based on the equilibrated homogeneous samples as detailed in the methods section.
Figure 3: (3(a)) Graphic representation of the system in a direct-coexistence simulation at the beginning of the calculation. (3(b)) Density profiles at the start of the simulation (dashed black curve) and at the end of the simulation, when the system has reached equilibrium (solid orange), for the $(AA)_{6}$ system at two different temperatures. (3(c)) Analogous results for the $(AB)_{6}$ systems for four different temperatures. In panels (3(b)) and (3(c)) characteristic snapshots of the final configurations are also shown as insets.
Figure 4: The phase diagram of the $(AB)_{6}$ system, displaying the density of the fluid, gas and liquid phases as extracted from direct-coexistence simulations. Note that the density of the gas at the lowest temperature ($10^{\circ}$ C) is approximated to zero, since no strand detached from the interfaces throughout the simulation. Note that the names ”gas phase” and ”liquid phase” are the ”implicit solvent” analog of polymer poor and polymer rich (network) phases.
We run very long simulations (up to $\approx 3\times 10^{10}$ simulation steps, corresponding to 2-3 months of simulation time on a NVIDIA A100 GPU), using the fraction of inter- and intra-molecular bonds as a proxy for equilibration, which we assume to be attained once these fractions reach a plateau value. We then compute density profiles along the elongated box, and classify each system as phase separated or homogeneous depending on whether the equilibrated profiles show the presence of interfaces or not.
Figure 3(b) shows the initial and final (equilibrium) density profiles for the $(AA)_{6}$ system, for which the direct coexistence simulations evolve toward an homogeneous configuration, providing strong evidence of the absence of phase separation in favor of a continuous cross-over from the dilute to the network state on increasing density. By contrast, Figure 3(c) shows that in the $(AB)_{6}$ case and at low temperature the density of the dense phase relaxes progressively to an equilibrium value, but in this case retaining a clear interface with a lower-density phase, indicating that a dilute solution of intra-bonded strands coexists with a network of mostly inter-bonded strands. It is only at $T=17.5^{\circ}$ C and above the $(AB)_{6}$ system approaches a homogeneous state for all densities.
From the value of the density in the dense and in the dilute phases at equilibrium it is possible to estimate the two coexisting densities. To do so in the phase-separated samples, we classify DNA strands as belonging to the interface, gas or liquid phase, making it possible to draw a phase diagram of the system, shown in Figure 4. As the temperature increases, the density of the coexisting phases approach each other, and from the absence of interfaces at $T=17.5^{\circ}$ C we can predict the critical temperature of the $(AB)_{6}$ system to lie between $15^{\circ}$ C and $17.5^{\circ}$ C. By evaluating the average between the two coexisting densities at $T=15^{\circ}$ C we can also provide a rough estimate for the critical density, $\rho_{c}\sigma^{3}=0.06$, corresponding to a strand concentration of $\approx 1$ mM or a weight density of $\approx 53$ mg/ml.
Figure 5: Normalised histograms of the bonds found in the phase-separating $(AB)_{6}$ system simulated at $T=15^{\circ}$ C for strands belonging to the gas and liquid phases (see text for definitions). Panels (a), (b) and (c) show the one-dimensional histograms for the intra-molecular, inter-molecular, and total bonds, respectively. Panels (d) and (e) show the two-dimensional histograms for the intra- and inter-molecular bonds of the strands found in the gas and liquid phases, respectively.
As shown in Fig. 4, for $T=15^{\circ}$ C the fraction of DNA strands in the gas is not negligible. This makes it possible to compare the properties of the single strands that make up the two phases for this particular state point. In particular, we investigate the difference in binding properties of the two phases. To this end, we show in Figure 5 the histograms of the inter-strand, intra-strand and total sticky sequences involved in bonds. The zig-zag pattern in the intra-strand histogram originates from the fact that each intra-strand bond involves two sticky sequences. As expected, the low density phase is rich in intra-molecular bonds and poor in inter-molecular bonds, while the reverse is true for the high density phase. At this temperature, which is the only one for which we have a sufficient number of gas strands to analyse, the two phases slightly differ in the total number of bonds ($\approx 9$ and $\approx 7.5$ average bonded stickers for the liquid and gas strands, respectively), signaling a weak energetic contribution to the difference in chemical potential. This small difference in the number of bonds will vanish at even lower temperatures, strongly suggesting that the two phases will have the same energy and thus that the separation is indeed driven by a strong entropic component stemming from the different conformations and different topology of the bonds adopted by the strands in the two phases.
Figure 6: (a) The average end-to-end distance, $r_{ee}$ as a function of the number of stickers involved in inter-molecular bonds, $N_{\rm inter}$, for the $(AA)_{6}$ and $(AB)_{6}$ homogeneous systems simulated at $T=15^{\circ}$ C and $\rho\sigma^{3}=0.16$. (b) $r_{ee}$ as a function of $N_{\rm inter}$ for the gas and liquid phases found in the direct-coexistence $(AB)_{6}$ system simulated at $T=15^{\circ}$ C. In both panels error bars show the mean-squared error, and therefore are absent for points obtained from single measurements (see e.g. the $N_{\rm inter}=10$ point in (b)).
To support this claim, we investigate the conformation of the strands extracted from the homogeneous simulations. Figure 6(a) shows a comparison between the average end-to-end distance of $(AA)_{6}$ and $(AB)_{6}$ strands, $r_{ee}$, as a function of the number of stickers involved in inter-molecular bonds. Since both the systems investigated are almost fully bonded, more inter-molecular bonds means fewer intra-molecular bonds, which tend to compact the strands. As a result, we observe that $r_{ee}$ increases monotonically with $N_{\rm inter}$. The roughly linear behaviour can be explained by assuming that each additional broken intra-molecular bond contributes to the end-to-end distance by the same amount, since each bond constraints portion of the strands that are approximately the same length. As expected, $(AB)_{6}$ strands are always smaller than $(AA)_{6}$ strands, since each intra-molecular bond closes a loop that is at least twice in size that of an $(AA)_{6}$ strand. It is this entropic penalty that shifts the balance towards inter-molecular bonds, in line with previous results on bead-spring models [20].
If we now focus on the phase-separated $(AB)_{6}$ system simulated at $T=15^{\circ}$ C, we see in Figure 6(b) that, as expected, in the gas the small number of inter-molecular bonds leads to compact conformations, whereas the liquid, with its higher inter-molecular connectivity, features larger end-to-end distances. Such a large difference in the conformational freedom of the strands tips the balance towards phase separation in the $(AB)_{6}$ system, providing a realistic example of the fully-entropic phase separation observed in bead-spring models.
IV Conclusions
In this work, we have presented a computational study of DNA-based associative polymers, designed to emulate the behavior of idealized APs, while at the same time providing an experimentally realisable system. The capacity to design sticky-end DNA sequences that are simultaneously self-complementary and mutually orthogonal makes this fully DNA-based associative polymer a versatile and powerful platform for encoding diverse bond architectures.
By leveraging the oxDNA model, we simulated single-stranded DNA constructs featuring either one or two types of self-complementary sticky sequences, separated by flexible poly-T spacers. Our findings show that at low temperatures on increasing concentration, systems with a single sticker type progressively form highly bonded, homogeneous networks, while those with alternating sticker types undergo a sharp first order phase separation, nucleating a connected network of polymers.
This phase separation arises from an entropic driving force associated with the suppression of intra-molecular bonding in favor of more flexible inter-molecular connections. The presence of phase separation in strands decorated with two types of stickers mirrors predictions from coarse-grained bead-spring models [20] and demonstrates that similar principles can be realized in realistic, sequence-programmable systems.
The results here presented not only validate previous theoretical work but also highlight the utility of DNA as a tunable platform for investigating the thermodynamics of associative networks. Our approach opens the door to the experimental realization of entropy-driven phase behavior in all-DNA systems, with potential implications for the design of responsive materials and hopefully for a deeper understanding of systems where phase separation of biomacromolecules is important [9, 26].